How to predict the number of bonds formation of P2O74- ? Please Explain It.

Grade:12th pass

2 Answers

Umakant biswal
7 years ago
@ styam
To predict the no of bond formation between p2o7 there is a formula
no of bond = full valence shell – no of valence electron
for example – hydrogen typically makes 1 bond beacuse its full valence shell is 2 , and the valence no is 1
so, in the above cases
in the case of p full valence shell is 8 , and in the case of o also the full valence shell is 8 .
so, full valence shell = 2(8)+7(8) = 16+ 56 = 72 -4 =68 ,
now calculate the valence electron , and minus it from 68 , u will get the answer .
HOPE IT CLEARS YOUR DOUBT
ALL THE BEST ,
IF NOT ABLE TO DO , THEN LET ME KNOW , I WILL SOLVE IT FULLY .
Satyam Anand
7 years ago
does this happen only with phosphorus or also with other elemnts of period 3 ?